OpenStudy (anonymous):

Create a scenario for a geometric sequence. For example, Anthony goes to the gym for __20____ minutes on Monday. Every day he ___increases______his gym time by ____10________. If he continues this pattern, how many minutes will he spend at the gym on the 5th day? Be sure to fill in the blanks with the words that will create a geometric sequence. Use your scenario to write the formula for the 5th term in your sequence, using sequence notation. Monday - 20 minutes Tuesday - 30 minutes Wednesday - 40 minutes Thursday - 50 minutes Friday - 60 minutes

OpenStudy (anonymous):

part 3 Use your scenario from part 2 to write a question that will lead to using the geometric series formula. Use the formula to solve for Sn in your scenario.

OpenStudy (anonymous):

I already have my data, I just need help to write a formula for my 5th term , using sequence notation and then i need help with part 3.

OpenStudy (ranga):

Give us your data.

OpenStudy (anonymous):

Anthony goes to the gym for __20____ minutes on Monday. Every day he ___increases______his gym time by ____10________. If he continues this pattern, how many minutes will he spend at the gym on the 5th day? Be sure to fill in the blanks with the words that will create a geometric sequence. Use your scenario to write the formula for the 5th term in your sequence, using sequence notation. Monday - 20 minutes Tuesday - 30 minutes Wednesday - 40 minutes Thursday - 50 minutes Friday - 60 minutes

OpenStudy (ranga):

That is an arithmetic sequence. Not a geometric sequnce.

OpenStudy (anonymous):

@ranga same here.

OpenStudy (ranga):

You can say he goes to the gym for 20 minutes on Monday and everyday he increases his gym time by 10%. Then calculate the time for Tue, Wed, etc.

OpenStudy (anonymous):

and i would keep multiyplying by 10 for each day ?

OpenStudy (ranga):

Not by 10. It is an increase of 10% (PERCENT). So you multiply by 1.1 each time (multiplying by 1.1 reflects an increase by 10%).

OpenStudy (ranga):

10% is 0.1

OpenStudy (anonymous):

so i do 20 times what?

OpenStudy (ranga):

20 * 1.1 = 22 minutes

OpenStudy (ranga):

Then 22 * 1.1 = 24.2 ....

OpenStudy (anonymous):

oh ok, so what about part 3?

OpenStudy (ranga):

Complete Thursday and Friday too.

OpenStudy (anonymous):

yeah i already wrote it part 3 Use your scenario from part 2 to write a question that will lead to using the geometric series formula. Use the formula to solve for Sn in your scenario.

OpenStudy (anonymous):

by the way the 'n' in 'Sn' is an exponent but at the bottom of S

OpenStudy (ranga):

"series" implies summing up or adding up. So if were were to ask the question how much time did Anthony spend at the gym from Monday to Friday then we will have to add the time from Monday to Friday..

OpenStudy (anonymous):

ok... but whats Sn

OpenStudy (anonymous):

whats the formula template im suppose to use to plug in whatever im suppose to plug in

OpenStudy (ranga):

There is a standard formula for the sum of the first n terms of a geometric series:\[\Large S _{n} = a \frac{ 1 - r^n }{ 1-r }\]

OpenStudy (ranga):

where a is the first term and r is the common ratio. Here a = 20 minutes and r = 1.1 and to find out the total time at the gym from Mon-Fri, n = 5 days. Plug the numbers into the formula and tell me what u get.

OpenStudy (anonymous):

wait so the total time from mon-fri is all the minutes added right ?

OpenStudy (ranga):

Yes. But they are asking you to use the formula to find the sum. Do that first and later also manually add up the times for the 5 days and see if they match.

OpenStudy (anonymous):

S(5)=20 1-1.1^5 / 1-1.1

OpenStudy (ranga):

yes. use calculator to find sum. make use of parenthesis properly.

OpenStudy (ranga):

no. 20 * ( 1 - 1.1^5 ) / ( 1 - 1.1) = ?

OpenStudy (ranga):

123 is not correct. Use parenthesis.

OpenStudy (anonymous):

122 ?

OpenStudy (ranga):

No. Make use of the parenthesis in your calculator and type them just as I wrote them above: 20 * ( 1 - 1.1^5 ) / ( 1 - 1.1) =

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

i got 122.102, u try it

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Yes. 122.102. Now list the gym time for the 5 days we did earlier. We did 3 days: 20, 22, 24.2 and we need data for Thu & Fri. List them here and using calculator add the 5 numbers to see if the total agrees with the total calculated using formula.

OpenStudy (anonymous):

Monday - 20 minutes Tuesday - 22 minutes Wednesday - 24 minutes Thursday - 27 minutes Friday - 30 minutes = 123

OpenStudy (ranga):

No. Mon - 20. Tue - 20 * 1.1 = 22, Wed - 22 * 1.1 = 24.2 Thu - 24.2 * 1.1 =

OpenStudy (anonymous):

27

OpenStudy (ranga):

Come on. You just have to use your calculator to multiply two numbers.

OpenStudy (ranga):

Thu - 26.62 minutes Fri - 26.62 * 1.1 = ?

OpenStudy (anonymous):

29.282

OpenStudy (anonymous):

rounded to 30

OpenStudy (ranga):

You can do the rounding off normally but here the sequence has to be geometric. That means the ratio must be the same when you divide successive numbers. if you round off the ratio will not be the same.

OpenStudy (anonymous):

ok wat now

OpenStudy (anonymous):

@tkhunny

OpenStudy (ranga):

These are the 5 numbers that you should list for part 2: Mon - 20 minutes Tue - 22 minutes Wed - 24.2 minutes Thu - 26.62 minutes Fri - 29.282 minutes And for your own verification, you can add these 5 numbers and see if it matches the total calculated using the formula earlier.

OpenStudy (tkhunny):

It is only a coincidence that we both picked 10% increases. There is nothing magic about that. You could use 20% increases and multiply by 1.2 each time.

OpenStudy (anonymous):

@ranga ... it does

OpenStudy (anonymous):

@ranga so now what ... they;re both the same so...

OpenStudy (ranga):

so you can be sure your calculations are correct. So all three parts have been answered.

OpenStudy (anonymous):

ok so this is what im submitting Task 3 Create a scenario for an arithmetic sequence. For example, Jasmine practices the piano for _10_____ minutes on Monday. Every day she ____increases_______ her practice time by __15 minutes_______. If she continues this pattern, how many minutes will she practice on the 7th day? Be sure to fill in the blanks with the words that will create an arithmetic sequence. Use your scenario to write the function for the 7th term in your sequence using sequence notation. Monday- 10 minutes Tuesday - 25 minutes Wednesday - 40 minutes Thursday - 55 minutes Friday - 70 minutes Saturday - 85 minutes Sunday - 100 minutes She will practice piano for 100 minutes on the 7th day, Sunday. function : An = a1 + d(n – 1) Day N = 10 +15(n-1) Day 7= 10+15 (7-1) part 2 Create a scenario for a geometric sequence. For example, Anthony goes to the gym for __20____ minutes on Monday. Every day he ___increases______his gym time by ____10%________. If he continues this pattern, how many minutes will he spend at the gym on the 5th day? Be sure to fill in the blanks with the words that will create a geometric sequence. Use your scenario to write the formula for the 5th term in your sequence, using sequence notation. (i rounded the decimals here) Mon - 20 minutes Tue - 22 minutes Wed - 24.2 minutes Thu - 26.62 minutes Fri - 29.282 minutes He will spend 29.282 minutes at the gym on the 5th day, Friday. part 3 Use your scenario from part 2 to write a question that will lead to using the geometric series formula. Use the formula to solve for Sn in your scenario. Sn=a*1−r^n/1−r where a is the first term and r is the common ratio. Here a = 20 minutes and r = 1.1 and to find out the total time at the gym from Mon-Fri, n = 5 days. S(5)=20 1-1.1^5 / 1-1.1 =122.102